OpenStudy (anonymous):

Five more than a number tripled is half the number. What is the number? A. 6 B. –2 C. –1 D. 1

OpenStudy (anonymous):

3x + 5 = x/2 solve or substitute and compare

OpenStudy (anonymous):

help??

OpenStudy (anonymous):

@douglaswinslowcooper It says tripled that's an exponent not multiplication.

OpenStudy (anonymous):

So it's x^3 + 5 = x/2